OpenStudy (anonymous):

How would you find the distance between two points if the line segment that them were vertical?

OpenStudy (anonymous):

Multiple choice

OpenStudy (uri):

Options?

OpenStudy (phi):

you would use the distance formula. Of course it is simple if the line goes straight up and down

OpenStudy (anonymous):

A.subtract the x-cordinate of the points and take the absolute value

OpenStudy (uri):

I have only learnt how to find distance with the distance formula so far.

OpenStudy (uri):

I think subract y-cordinate

OpenStudy (anonymous):

B.Add the x-cordinate of the points and take the absolute value

OpenStudy (anonymous):

C.Add the y-cordinate of the points and take the absolute value

OpenStudy (anonymous):

You think?

OpenStudy (uri):

Subract the y-cordinate...do you have that option?

OpenStudy (anonymous):

yea thats D.

OpenStudy (uri):

Hm

OpenStudy (phi):

if the line goes up and down, all the points have the same x value the distance formula is \[ \sqrt{ (x_1 - x_2)^2 + (y_1- y_2)^2 } \] if the x values are all the same, then \( x_1 - x_2= 0\) and the formula simplifies to \[ \sqrt{(y_1- y_2)^2 } \] now take the square root \[ (y_1-y_2) \] however, we only take the positive square root, so we should show this by absolute value \[ |y_1-y_2| \]